OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 13m63. The length of its base is twice the width. material for the base costs $9 per square meter. Material for the sides costs $13 per square meter. Find the cost of materials for the cheapest such container.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

first step is to build a cost equation: b*9+s*13=totalcost equation for base is length*width or b=lw sides = perimeter * height and perimeter can be found with (2l+2w) or 2lengths + 2 heights and l is equal to 2w so that can be subbed in where ever that appears b=2w*w or b=2w^2 sides = 2(2w) + 2w *h and the whole thing b*h=13m^3.......I think that is all the equations then find first and second derivative of total cost to find min cost

OpenStudy (anonymous):

why don't i read those before I hit post (2l+2w) or 2lengths + 2widths (not heights)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

just had to work that one myself recently